Can you solve the linear system by using the substitution method and showing details please. Thank you
3x+4y=-3
2x+y=8

Answer by scott8148(6628)

solving the 2nd eqn for y ___ y = 8 - 2x

substituting the value for y into the 1st eqn ___ 3x + 4(8 - 2x) = -3 ___ 3x + 32 - 8x = -3 ___ -5x = -35

solve for x, and then substitute back to find y